# NODEC: Neural ODE For Optimal Control of Unknown Dynamical Systems

### 03 Jan 2024 v1

@@ Results
Figure (5) indicates the full training process of NODEC for controlling the CartPole system with four pairs of plots. The pair on the top corresponds to the initial training alternation ($k=0$), while the bottom represents the fully trained model after $K=12$ alternations. The pairs in the middle correspond to intermediate training alternations. In each plot pair, the left plot illustrates the trajectory roll-outs under the real dynamics $f$ (shown in dashed lines) and under the current dynamics learner ${g}_{\gamma}^{k}$ using the controller. Since the states in this case have four dimensions, we represent the trajectory roll-outs by plotting the values of each state dimension over time, rather than directly visualizing the trajectories in the state space. It is important to note that we visualize the results for a single initial state from the batch of data. The right plot in each pair displays the control signals generated along the trajectories.
We can observe that NODEC controller $h_{\theta}^{K}$ successfully controls the CartPole system after K alternations, as it keeps the system evolution near the target state (0,0,0,0) which corresponds to the upright position of the pole, and the trajectory roll-outs in learned dynamics are close to real dynamics. Again, the training losses for dynamics learner and controller correspond to these four alternations are shown in Figure (9) Appendix C.
One thing to note here is the data efficiency of NODEC algorithm. As indicated in Algorithm 1 line 5, we only need to interact with the real environment and collect real trajectories at each alternation $k$ for each initial state in our batch. Therefore, we only need to collect in total $K*N$ trajectories or episodes from the real environment where $N$ is the batch size. For this particular experiment, we only collected less than 60 trajectories before successfully solving the CartPole tasks, which is more data-efficient compared to RL methods that uses hundreds or thousands of trajectories.
@@ 4 Conclusion
We introduced a novel framework denoted as NODEC that consists of a coupled neural ODE for controlling dynamical systems with unknown dynamics. This is achieved via an alternative training between two neural networks in our NODEC model: the dynamics learner and the controller. Through an intriguing interplay and mutual supervision between these two neural networks, the model learns the optimal control as well as the system dynamics. We applied NODEC in different dynamical systems and demonstrated the effectiveness of NODEC for learning the optimal control with high data efficiency.
This manuscript presents a detailed and technical exposition of the method introduced in Chapter 3 (’TEL for Optimal Control: NODEC’) of the thesis titled ’Theory Embedded Learning’ [20]. For further reference, the relevant code implementations can be accessed here
@@ Appendix A Alternative training motivation
Figure 6 below depicts the trajectory roll-outs under ${g}{\gamma}(x,{h}{\theta}(x,t),t)$ using already well-trained ${g}{\gamma}$ and ${h}{\theta}$ (after K alternations in Figure 3. The figure also includes visualizations of the true vector field of $f$ and the vector field of the learned dynamics ${g}$. While the vector field approximation may not be ideal across most of the state space, it fits perfectly in the vicinity of the trajectory roll-outs (as indicated by the black circles where dark red and dark blue vector fields are aligned). This demonstrate the points we made in section 2.4: we do not need a perfect fit dynamics learner, but rather just a good local approximation of the dynamics near the state space where the current controllers can lead to. And this local good approximation already suffices for learning the optimal controls.
Figure 6: Vector field visualization of trained NODEC for Ax + Bu task
@@ Appendix B Compare with analytical optimal control
Here we compare the trained NODEC controller in section 3 Figure 3 with the analytical optimal controller given by Equation 7. The left plot in Figure 7 above shows the trajectory roll-out in learned dynamics (which is almost the same as the real dynamics as indicated by the last pair of plots in Figure 3). The black trajectories are roll-out using trained controller with NODEC (the same trajectories in the last pair in Figure 3), while the blue trajectories are roll-out using optimal controllers. The high similarity between trajectory roll-outs as well as control signals indicate that through alternative training, our NODEC model can indeed learn optimal controls without any prior knowledge about system dynamics.
Figure 7: Comparison of NODEC controller with optimal controller

### 04 Feb 2024 v3

@@ Results
Figure (5) indicates the full training process of NC for controlling the CartPole system with four pairs of plots. The pair on the top corresponds to the initial training alternation ($k=0$), while the bottom represents the fully trained model after $K=12$ alternations. The pairs in the middle correspond to intermediate training alternations. In each plot pair, the left plot illustrates the trajectory roll-outs under the real dynamics $f$ (shown in dashed lines) and under the current dynamics learner ${g}_{\gamma}^{k}$ using the controller. Since the states in this case have four dimensions, we represent the trajectory roll-outs by plotting the values of each state dimension over time, rather than directly visualizing the trajectories in the state space. It is important to note that we visualize the results for a single initial state from the batch of data. The right plot in each pair displays the control signals generated along the trajectories.
We can observe that NC controller $h_{\theta}^{K}$ successfully controls the CartPole system after K alternations, as it keeps the system evolution near the target state (0,0,0,0) which corresponds to the upright position of the pole, and the trajectory roll-outs in learned dynamics are close to real dynamics. Again, the training losses for dynamics learner and controller correspond to these four alternations are shown in Figure (9) Appendix C.
One thing to note here is the data efficiency of NC algorithm. As indicated in Algorithm 1 line 5, we only need to interact with the real environment and collect real trajectories at each alternation $k$ for each initial state in our batch. Therefore, we only need to collect in total $K*N$ trajectories or episodes from the real environment where $N$ is the batch size. For this particular experiment, we only collected less than 60 trajectories before successfully solving the CartPole tasks, which is more data-efficient compared to RL methods that uses hundreds or thousands of trajectories.
@@ 4 Conclusion
We introduced a novel framework denoted as NC that consists of a coupled neural ODE for controlling dynamical systems with unknown dynamics. This is achieved via an alternative training between two neural networks in our NC model: the dynamics learner and the controller. Through an intriguing interplay and mutual supervision between these two neural networks, the model learns the optimal control as well as the system dynamics. We applied NC in different dynamical systems and demonstrated the effectiveness of NC for learning the optimal control with high data efficiency.
This manuscript presents a detailed and technical exposition of the method introduced in Chapter 3 (’TEL for Optimal Control: NC’) of the thesis titled ’Theory Embedded Learning’ [28]. The relevant code implementations can be accessed here.
@@ Appendix A Alternative training motivation
Figure 6 below depicts the trajectory roll-outs under ${g}{\gamma}(x,{h}{\theta}(x,t),t)$ using already well-trained ${g}{\gamma}$ and ${h}{\theta}$ (after K alternations in Figure 3. The figure also includes visualizations of the true vector field of $f$ and the vector field of the learned dynamics ${g}$. While the vector field approximation may not be ideal across most of the state space, it fits perfectly in the vicinity of the trajectory roll-outs (as indicated by the black circles where dark red and dark blue vector fields are aligned). This demonstrate the points we made in section 2.3.3: we do not need a perfect fit dynamics learner, but rather just a good local approximation of the dynamics near the state space where the current controllers can lead to. And this local good approximation already suffices for learning the optimal controls.
Figure 6: Vector field visualization of trained NC for Ax + Bu task
@@ Appendix B Compare with analytical optimal control
Here we compare the trained NC controller in section 3 Figure 3 with the analytical optimal controller given by Equation 7. The left plot in Figure 7 above shows the trajectory roll-out in learned dynamics (which is almost the same as the real dynamics as indicated by the last pair of plots in Figure 3). The black trajectories are roll-out using trained controller with NC (the same trajectories in the last pair in Figure 3), while the blue trajectories are roll-out using optimal controllers. The high similarity between trajectory roll-outs as well as control signals indicate that through alternative training, our NC model can indeed learn optimal controls without any prior knowledge about system dynamics.
Figure 7: Comparison of NC controller with optimal controller